# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v1

@@ 3.2 Stepsize bounds
where the quantity on the right-hand side is at most zero. Together with the Lipschitz continuity of $\nabla f$ and (3.37) with $L=L_{{}_{\rm{f}}}$, we have for any $\alpha\in(0,\hat{\alpha}]$ that
$V_{k-1}-V_{k}$ $\geq\left(-\frac{\hat{\alpha}}{\tau}+\frac{1}{2(L_{{}_{\rm{f}}}-\rho_{{}_{\rm{f}}})}\right)L_{{}_{\rm{f}}}^{2}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}$
$\quad+\left[\frac{2-\tau}{2\tau\hat{\alpha}}-\frac{L_{{}_{\rm{h}}}}{2}-\frac{\rho_{{}_{\rm{f}}}L_{{}_{\rm{f}}}}{2(L_{{}_{\rm{f}}}-\rho_{{}_{\rm{f}}})}+\frac{\tau-1}{\tau}L_{{}_{\rm{h}}}-\frac{\hat{\alpha}}{\tau}L_{{}_{\rm{h}}}L_{{}_{\rm{f}}}\right]{\left\|{x^{k}-x^{k-1}}\right\|}^{2}$
@@
$=\left(\frac{2-\tau}{2\tau\hat{\alpha}}-\hat{\alpha}\frac{L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{\tau}+\frac{L_{{}_{\rm{f}}}}{2}-\frac{L_{{}_{\rm{h}}}}{2}+\frac{\tau-1}{\tau}L_{{}_{\rm{h}}}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}$
$=-\frac{c(\hat{\alpha})}{2\tau\hat{\alpha}}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2},$ (3.38)
@@
$c(\hat{\alpha})$ $\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\hat{\alpha}-(2-\tau)$
$=\left(\hat{\alpha}-\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)\left(2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}+(2-\tau)(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\right),$ (3.39)
which is nonpositive when $\hat{\alpha}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. Hence, (3.38) holds with nonpositive $c(\hat{\alpha})$ when $\hat{\alpha}\in\left[\frac{\tau}{2(L_{{}_{\rm{f}}}-\rho_{{}_{\rm{f}}})},\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right]$, which is a nonempty interval due to our hypothesis that $(2-\tau)L_{{}_{\rm{f}}}-2\rho_{{}_{\rm{f}}}\geq\tau L_{{}_{\rm{h}}}$. In particular, we may choose the maximum allowable stepsize $\hat{\alpha}=\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. Note that it is clear that $c(\alpha)<0$ when $\alpha<\bar{\alpha}$.
@@
$V_{k-1}-V_{k}$ $\geq\left[\frac{2-\tau}{2\tau\hat{\alpha}(L)}-\frac{L_{{}_{\rm{h}}}}{2}-\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}+\frac{\tau-1}{\tau}L_{{}_{\rm{h}}}-\frac{\hat{\alpha}(L)}{\tau}L_{{}_{\rm{h}}}L_{{}_{\rm{f}}}\right]{\left\|{x^{k}-x^{k-1}}\right\|}^{2}$
$\qquad+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2}$
$=-\frac{c(\hat{\alpha}(L))}{2\tau\hat{\alpha}(L)}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2},$
@@
Since (3.40) implies that $\hat{\alpha}(L_{{}_{\rm{f}}})>\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$, it follows that $c(\hat{\alpha}(L_{{}_{\rm{f}}}))>0$. Hence, $L^{*}\neq L_{{}_{\rm{f}}}$ by the definition of $L^{*}$ in (3.42). The claim that $L^{*}=\eta^{*}+\rho_{{}_{\rm{f}}}$ now follows. We also note that since $c(\hat{\alpha}(L^{*}))\leq 0$, it follows from (3.41) that $c(\alpha)<0$ for any $\alpha<\hat{\alpha}(L^{*})$, where $c(\alpha)$ is as defined in (3.27).
@@
To gain insight on the magnitude of the stepsize upper bound $\bar{\alpha}=\frac{\tau}{2\eta^{*}}$ in the second case of the above proof, consider $L\coloneqq\frac{\tau L_{{}_{\rm{h}}}+2\rho_{{}_{\rm{f}}}}{2-\tau}$. By (3.40), $L>L_{{}_{\rm{f}}}$ and $L>\rho_{{}_{\rm{f}}}$. In addition, for this particular choice of $L$, $c(\hat{\alpha}(L))\leq 0$. By (3.42), it holds that $L\geq L^{*}$, ensuring $\frac{\tau}{2\eta^{*}}\geq\frac{\tau}{2(L-\rho_{{}_{\rm{f}}})}=\frac{2-\tau}{2(L_{{}_{\rm{h}}}+\rho_{{}_{\rm{f}}})}$. To obtain an upper bound, we note that from (3.43), it can be verified that an alternative way to express $c(\hat{\alpha}(L))$ is
@@
Note that the last two terms are nonnegative when $L=L^{*}$ since $L^{*}\geq L_{{}_{\rm{f}}}$ and $L^{*}>\rho_{{}_{\rm{f}}}$. Since $c(\hat{\alpha}(L^{*}))=0$, it follows that $\hat{\alpha}(L^{*})-\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\leq 0$. That is, we have $\frac{\tau}{2\eta^{*}}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. In summary, we have the estimate $\frac{2-\tau}{2(L_{{}_{\rm{h}}}+\rho_{{}_{\rm{f}}})}\leq\frac{\tau}{2\eta^{*}}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$ and therefore
$\bar{\alpha}\leq\frac{1}{L_{f}+L_{h}}$ (3.44)
when $\tau\in(0,1]$.
@@
Given any $\tau\geq 1$, we have from (3.35) and the monotonicity of $\nabla h-\sigma_{{}_{\rm{h}}}Id$ that
@@
The rest of the proof proceeds using arguments similar to those in the proof of Theorem 3.5. ∎
@@
Similar to Remark 3.6, the bound Eq. 3.44 also holds for $\tau\in(1,2)$. To see this, note that $c(\alpha)$ given in (3.45) can also be written as
@@
where $\hat{c}(\alpha)\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\alpha-(2-\tau)$. Since $\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$ is the positive root of $\hat{c}(\alpha)$ (see (3.39)) and $L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}}\geq 0$ by Eq. 2.5, it holds that $c\left(\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)=\frac{2(\tau-1)(L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}})}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\geq 0$. Since $c(0)=\tau-2<0$ and $\bar{\alpha}_{1}$ is the positive root of $c(\alpha)$, it follows that $\bar{\alpha}_{1}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. On the other hand, following the same arguments in Remark 3.6, it can be shown that $\tau/(2\eta^{*})\leq\bar{\alpha}_{1}$, so Eq. 3.44 holds as claimed.
@@
for any $x,\hat{x}\in{\rm I}\!{\rm R}^{n}$ and any $\mu\in[0,1]$. Following the arguments in Lemma 3.4, we obtain
$V_{\Lambda,\xi}(y,z,x)\geq$ $~{}Q_{\alpha(f+h)}(y^{\scriptscriptstyle+};\hat{x})+p(y^{\scriptscriptstyle+})+g(y^{\scriptscriptstyle+})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle$
@@
By using this bound and inequality (3.46) with $\sigma_{{}_{\rm{h}}}=-\rho_{{}_{\rm{h}}}$, we obtain by following the same arguments in the proof of Theorem 3.5 that
@@
$\left(-\frac{\alpha}{\tau}+\frac{\mu}{2L_{{}_{\rm{f}}}}\right){\left\|{\nabla f(x^{k-1})-\nabla f(x^{k})}\right\|}^{2}$ $=\frac{\mu L_{{}_{\rm{h}}}}{2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{\left\|{\nabla f(x^{k-1})-\nabla f(x^{k})}\right\|}^{2}$
$\geq\frac{\mu L_{{}_{\rm{h}}}\sigma_{{}_{\rm{f}}}^{2}}{2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{\left\|{x^{k-1}-x^{k}}\right\|}^{2},$
where the inequality follows from the strong convexity of $f$. Therefore,
$V_{k-1}-V_{k}\geq-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2},$
where $c(\alpha)$ is given by (3.50). To obtain a nonincreasing sequence $\{V_{k}\}$, we need to find the range of $\mu$ that makes $c(\alpha)\leq 0$. Plugging in $\alpha=\frac{\tau\mu}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$ in (3.50), some algebraic calculations lead to
@@
That is to say, the first coordinate of the vertex of the parabola defined by $r(\mu)$ lies in $(0,\frac{1}{2}]$. Since $\tau\geq 2$, it follows that $0\leq\mu_{*}\leq\mu^{*}\leq 1$. Therefore, $c(\alpha)\leq 0$ for $\alpha=\frac{\tau\mu}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$ with $\mu\in[\mu_{*},\mu^{*}]$, and $c(\alpha)<0$ if $\mu\in(\mu_{*},\mu^{*})$. This completes the proof. ∎
@@
Suppose that $\tau=2$. Then Eqs. 3.47 and 3.48 are equivalent to having $\sigma_{{}_{\rm{f}}}>L_{{}_{\rm{h}}}+\rho_{{}_{\rm{h}}}$, and the roots of (3.49) are $\mu_{*}=0$ and $\mu^{*}=\frac{L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})(\sigma_{{}_{\rm{f}}}-L_{{}_{\rm{h}}}-\rho_{{}_{\rm{h}}})}{L_{{}_{\rm{h}}}(L_{{}_{\rm{f}}}^{2}-\sigma_{{}_{\rm{f}}}^{2})+\sigma_{{}_{\rm{f}}}L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$. Hence, $c(\alpha)$ is strictly negative for all $\alpha$ such that
@@
If $h=0$, the condition (3.48) automatically holds since $\sigma_{{}_{\rm{f}}}>0$. Moreover, the condition (3.47) with the constraint $\tau\geq 2$ is equivalent to having
@@
$\frac{\tau}{2L_{{}_{\rm{f}}}}\left(\frac{1}{2}-\frac{\sqrt{\nu(\nu\tau^{2}-8\tau+16)}}{2\tau\nu}\right)<\alpha<\frac{\tau}{2L_{{}_{\rm{f}}}}\left(\frac{1}{2}+\frac{\sqrt{\nu(\nu\tau^{2}-8\tau+16)}}{2\tau\nu}\right),$ (3.52)
or equivalently,
$\frac{\tau\nu-\sqrt{\nu(\nu\tau^{2}-8\tau+16)}}{4\sigma_{{}_{\rm{f}}}}<\alpha<\frac{\tau\nu+\sqrt{\nu(\nu\tau^{2}-8\tau+16)}}{4\sigma_{{}_{\rm{f}}}},$
which are the bounds obtained in (Themelis and Patrinos, 2020, Theorem 4.1), where the case $h=p=0$ is considered. However, in the said result, the analysis restricts $\tau$ to satisfy only the first condition in (3.51), due to their imposed constraint that $\alpha L_{{}_{\rm{f}}}$ must be at most 1. Meanwhile, the analysis we provide in the proof of Theorem 3.12 does not require this condition to establish the nonincreasing property, and therefore we have shown that the range of $\tau$ can be widened to include those that satisfy $\tau\geq\frac{4}{1-\sqrt{1-\nu}}$. For instance, if we are given $\nu=3/4$, following Eq. 3.51 we can choose $\tau=12$. By (3.52), for this value we may allow any stepsize $\alpha$ that satisfies $3-\frac{\sqrt{21}}{3}<\alpha L_{{}_{\rm{f}}}<3+\frac{\sqrt{21}}{3}$, and the lower bound is strictly greater than 1. Nevertheless, we point out that this wider range of $\tau$ provided above is only sufficient to guarantee monotonicity of $\{V_{k}\}$. If we want to establish boundedness of a sequence generated by Algorithm 1, the restriction that $\alpha(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})<1$ will inevitably be required (see Proposition 3.15).

### v2

@@ 3.2 Stepsize bounds
where the quantity on the right-hand side is at most zero. Together with the Lipschitz continuity of $\nabla f$ and (3.37) with $L=L_{{}_{\rm{f}}}$, for any $\alpha\in(0,\hat{\alpha}]$ we have
$V_{k-1}-V_{k}$ $\geq\left(\frac{2-\tau}{2\tau\hat{\alpha}}-\hat{\alpha}\frac{L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{\tau}+\frac{L_{{}_{\rm{f}}}}{2}-\frac{L_{{}_{\rm{h}}}}{2}+\frac{\tau-1}{\tau}L_{{}_{\rm{h}}}\right){\left\|{x^{k}-x^{k-1}}\right\|}^{2}$
@@
$=-\frac{c(\hat{\alpha})}{2\tau\hat{\alpha}}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$
$\quad+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2},$ (3.38)
@@
$\begin{split}c(\hat{\alpha})&\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\hat{\alpha}-(2-\tau)\\ &=\left(\hat{\alpha}-\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)\left(2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\hat{\alpha}+(2-\tau)(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\right)\end{split},$ (3.39)
which is nonpositive when $\hat{\alpha}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. Hence, (3.38) holds with nonpositive $c(\hat{\alpha})$ when $\hat{\alpha}\in\left[\frac{\tau}{2(L_{{}_{\rm{f}}}-\rho_{{}_{\rm{f}}})},\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right]$, which is a nonempty interval due to our hypothesis that $(2-\tau)L_{{}_{\rm{f}}}-2\rho_{{}_{\rm{f}}}\geq\tau L_{{}_{\rm{h}}}$. It is clear that $c(\alpha)<0$ when $\alpha<\bar{\alpha}$.
@@
$V_{k-1}-V_{k}$ $\geq-\frac{c(\hat{\alpha}(L))}{2\tau\hat{\alpha}(L)}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$ (3.41)
$\quad+\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y^{k}-y^{k+1}}\right\|}^{2},$
@@
Since (3.40) implies that $\hat{\alpha}(L_{{}_{\rm{f}}})>\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$, it follows that $c(\hat{\alpha}(L_{{}_{\rm{f}}}))>0$. Hence, $L^{*}\neq L_{{}_{\rm{f}}}$ by the definition of $L^{*}$ in (3.43). The claim that $L^{*}=\eta^{*}+\rho_{{}_{\rm{f}}}$ now follows. We also note that since $c(\hat{\alpha}(L^{*}))\leq 0$, it follows from (3.42) that $c(\alpha)<0$ for any $\alpha<\hat{\alpha}(L^{*})$, where $c(\alpha)$ is as defined in (3.27). ∎
@@
To gain insight on the magnitude of the stepsize upper bound $\bar{\alpha}=\frac{\tau}{2\eta^{*}}$ in the second case of the above proof, consider $L\coloneqq\frac{\tau L_{{}_{\rm{h}}}+2\rho_{{}_{\rm{f}}}}{2-\tau}$. By (3.40), $L>L_{{}_{\rm{f}}}$ and $L>\rho_{{}_{\rm{f}}}$. In addition, for this choice of $L$, $c(\hat{\alpha}(L))\leq 0$. By (3.43), it holds that $L\geq L^{*}$, ensuring $\frac{\tau}{2\eta^{*}}\geq\frac{\tau}{2(L-\rho_{{}_{\rm{f}}})}=\frac{2-\tau}{2(L_{{}_{\rm{h}}}+\rho_{{}_{\rm{f}}})}$. To obtain an upper bound, we note that from (3.44), it can be verified that an alternative way to express $c(\hat{\alpha}(L))$ is
@@
Note that the last two terms are nonnegative when $L=L^{*}$ since $L^{*}\geq L_{{}_{\rm{f}}}$ and $L^{*}>\rho_{{}_{\rm{f}}}$. Since $c(\hat{\alpha}(L^{*}))=0$, it follows that $\hat{\alpha}(L^{*})-\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\leq 0$. That is, $\frac{\tau}{2\eta^{*}}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. In summary, when $\tau\in(0,1]$, we have $\frac{2-\tau}{2(L_{{}_{\rm{h}}}+\rho_{{}_{\rm{f}}})}\leq\frac{\tau}{2\eta^{*}}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$ and therefore
$\bar{\alpha}\leq\frac{1}{L_{f}+L_{h}}.$ (3.45)
@@
Given $\tau\geq 1$, we have from (3.35) and the monotonicity of $\nabla h-\sigma_{{}_{\rm{h}}}Id$ that
@@
The rest of the proof follows arguments similar to those in the proof of Theorem 3.5. ∎
@@
Similar to Remark 3.6, the bound Eq. 3.45 also holds for $\tau\in(1,2)$. To see this, note that $c(\alpha)$ in (3.46) can also be written as
@@
where $\hat{c}(\alpha)\coloneqq 2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})\alpha^{2}+\left((2-\tau)L_{{}_{\rm{h}}}-\tau L_{{}_{\rm{f}}}\right)\alpha-(2-\tau)$. Since $\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$ is the positive root of $\hat{c}(\alpha)$ (see (3.39)) and $L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}}\geq 0$ by Remark 2.2, it holds that $c\left(\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\right)=\frac{2(\tau-1)(L_{{}_{\rm{h}}}-\sigma_{{}_{\rm{h}}})}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\geq 0$. Since $c(0)=\tau-2<0$ and $\bar{\alpha}_{1}$ is the positive root of $c(\alpha)$, it follows that $\bar{\alpha}_{1}\leq\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}$. On the other hand, following the same arguments in Remark 3.6, it can be shown that $\tau/(2\eta^{*})\leq\bar{\alpha}_{1}$, so Eq. 3.45 holds as claimed.
@@
for any $x,\hat{x}\in{\rm I}\!{\rm R}^{n}$ and $\mu\in[0,1]$. Following the arguments in Lemma 3.4, we obtain
$V_{\Lambda,\xi}(y,z,x)\geq$
$\quad Q_{\alpha(f+h)}(y^{\scriptscriptstyle+};\hat{x})+p(y^{\scriptscriptstyle+})+g(y^{\scriptscriptstyle+})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle$
@@
By using this bound and inequality (3.47) with $\sigma_{{}_{\rm{h}}}=-\rho_{{}_{\rm{h}}}$, we obtain by following the same arguments and calculations in the proof of Theorem 3.5 that
@@
$\left(-\frac{\alpha}{\tau}+\frac{\mu}{2L_{{}_{\rm{f}}}}\right){\left\|{\nabla f(x^{k-1})-\nabla f(x^{k})}\right\|}^{2}\geq\frac{\mu L_{{}_{\rm{h}}}\sigma_{{}_{\rm{f}}}^{2}}{2L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}{\left\|{x^{k-1}-x^{k}}\right\|}^{2},$
where the inequality follows from the strong convexity of $f$. Therefore, (3.24) holds with $c(\alpha)$ given by (3.51). To obtain a nonincreasing sequence $\{V_{k}\}$, we need to find the range of $\mu$ that makes $c(\alpha)\leq 0$. Plugging in $\alpha=\frac{\tau\mu}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$ in (3.51), some algebraic calculations lead to
@@
That is to say, the first coordinate of the vertex of the parabola defined by $r(\mu)$ lies in $(0,\frac{1}{2}]$. Since $\tau\geq 2$, it follows that $0\leq\mu_{*}<\mu^{*}\leq 1$. Thus, $c(\alpha)\leq 0$ for $\alpha=\frac{\tau\mu}{2(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$ with $\mu\in[\mu_{*},\mu^{*}]$, and $c(\alpha)<0$ if $\mu\in(\mu_{*},\mu^{*})$. This completes the proof. ∎
@@
Suppose that $\tau=2$. Then Eqs. 3.48 and 3.49 are equivalent to having $\sigma_{{}_{\rm{f}}}>L_{{}_{\rm{h}}}+\rho_{{}_{\rm{h}}}$, and the roots of (3.50) are $\mu_{*}=0$ and $\mu^{*}=\frac{L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})(\sigma_{{}_{\rm{f}}}-L_{{}_{\rm{h}}}-\rho_{{}_{\rm{h}}})}{L_{{}_{\rm{h}}}(L_{{}_{\rm{f}}}^{2}-\sigma_{{}_{\rm{f}}}^{2})+\sigma_{{}_{\rm{f}}}L_{{}_{\rm{f}}}(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})}$. Hence, $c(\alpha)$ is strictly negative for all $\alpha$ such that
@@
If $h\equiv 0$, the condition (3.49) automatically holds since $\sigma_{{}_{\rm{f}}}>0$. Moreover, the condition (3.48) with the constraint $\tau\geq 2$ is equivalent to having
@@
$\frac{\tau\nu-\sqrt{\nu(\nu\tau^{2}-8\tau+16)}}{4\sigma_{{}_{\rm{f}}}}<\alpha<\frac{\tau\nu+\sqrt{\nu(\nu\tau^{2}-8\tau+16)}}{4\sigma_{{}_{\rm{f}}}},$ (3.53)
which are the bounds obtained in (Themelis and Patrinos, 2020, Theorem 4.1), where the case $h\equiv p\equiv 0$ is considered. However, in the said result, the analysis restricts $\tau$ to satisfy only the first condition in (3.52), due to their imposed constraint that $\alpha L_{{}_{\rm{f}}}$ must be at most 1. Meanwhile, the analysis we provide in the proof of Theorem 3.12 does not require this condition to establish the nonincreasing property, and therefore we have shown that the range of $\tau$ can be widened to include those that satisfy $\tau\geq\frac{4}{1-\sqrt{1-\nu}}$. For instance, if we are given $\nu=3/4$, following Eq. 3.52 we can choose $\tau=12$. By (3.53), for this value we may allow any stepsize $\alpha$ that satisfies $3-\frac{\sqrt{21}}{3}<\alpha L_{{}_{\rm{f}}}<3+\frac{\sqrt{21}}{3}$, and the lower bound is strictly greater than 1. Nevertheless, we point out that this wider range of $\tau$ provided above is only sufficient to guarantee monotonicity of $\{V_{k}\}$. If we want to establish boundedness of a sequence generated by Algorithm 1, the restriction that $\alpha(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})<1$ will inevitably be required (see Proposition 3.15).